If eight coins are thrown randomly, show that the probability of obtaining at least 5 heads is 93/256
Answer #1

Sample size , n =    8
Probability of an event of interest, p =   0.5000


Binomial probability is given by , P(X=x) = C(n,x)*px*(1-p)(n-x)      

p(at least 5 heads)= p(5 heads)+ p(6heads)+p(7 heads)+p(8 heads)

=0.2188+0.1094+ 0.0313+0.0039

= 0.3633